OpenStudy (anonymous):

Lori Bray drives a parcel delivery truck. Her usual route consists of 10 stops. How many ways can Lori travel to these stops and return to her warehouse? A. 90 B. 3,628,800 C. 362,880 D. 10

OpenStudy (unklerhaukus):

well the first delivery could be any of the 10 the second delivery could be any of the 9 remaining the next delivery could be any of the remaining, ... the second last delivery can be either of the two remaining the last delivery could only be the last remain one, hence the probability is 10x9x8x7x6x5x4x3x2x1 =10! =3628800

OpenStudy (unklerhaukus):

*probability is 1/10x9x8x7x6x5x4x3x2x1 =1/10 =1/3628800 there are 10!=3628800 possibilities